Diffusion Morning Report 08/03/07

Ted Berg tberg at stanford.edu
Fri Aug 3 06:27:28 PDT 2007

Good Morning All.
    Looking OK this morning.  POCl3 still under investigation. TEOS2 and 
BBr3 under observation. LTO going down on Monday for pull and clean.Burn 
box cleaned this morning.  No other major problems at this time.

  	Date: 	8/3/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	Comments 	  	  	 
Tylan 1 	up 	  	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	down 	tube clean 	  	  	 
Tylan 7-TEOS2 	yellow 	observation 	9.99/30 	  	 
Tylan 8-SiGe 	up 	  	19.63/30 	  	 
Tylan 9-Poly 	up 	  	16.97/30 	  	 
Tylan 10-Nitride 	up 	  	24.83/30 	  	 
Tylan 11-BPSG-LTO 	yellow 	particles? 	23.64/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	up 	  	  	  	 
Thermco Poly2 	future 	  	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
Burn Box 	Up 	  	   	  	 
Ebeam 	up 	  	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	A on line 600 psi 	 changeover working well  	  	 

-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070803/b2409320/attachment.html>

More information about the diffusion mailing list